OpenStudy (anonymous):

A baseball is thrown upwards from a height of 5 feet with an initial speed of 64 feet per second, and its height h (in feet) from the ground is given by h(t) = 5 + 64t – 16t2 as shown in the graph where t is time in seconds. Using a graphing calculator, at what time does the ball reach its maximum height? 0 1 2 3

OpenStudy (anonymous):

Is this 2 or C?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@amistre64

OpenStudy (phi):

\[ h(t) = – 16t^2 + 64t +5\] they ask you to use a graphing calculator. but you can use algebra. match your equation to \[ y = a x^2 + bx + c\] the peak occurs at x= -b/(2a) in your problem, a matches -16 and b matches +64 and the peak is at t= -b/(2a)= -64/(2*-16)

OpenStudy (anonymous):

thas 2

OpenStudy (phi):

yes, t=2 is the top